I am attempting to go over my 13 yo grandson's math homework and my age has humbled me to ask for help from you. I would greatly appreciate your assistance due to the late hour, I apologize but I promised I would check his work and I am lost.
A loving Nana in NC
Determine whether the given value is a solution of the equation. YES OR NO(Example 1)
1. 23 = x - 9; x = 14
2. n __ 13 = 4; n = 52
3. 14 + x = 46; x = 32
4. 17y = 85; y = 5
5. 25 = k _ 5 ; k = 5
6. 2.5n = 45; n = 18
7. 21 = m + 9; m = 11
8. 21 - h = 15; h = 6
9. d - 4 = 19; d = 15
10. 5 + x = 47; x = 52
11. w - 9 = 0; w = 9
12. 5q = 31; q = 13
13. Each floor of a hotel has r rooms. On 8 floors, there are a total of 256 rooms. Write an equation to represent this situation,


Answer by ikleyn(52915)

.
Determine whether the given value is a solution of the equation. YES OR NO(Example 1)

1. 23 = x - 9; x = 14          NO   (14 - 9 = 5, not 23)

2.   n __ 13   = 4; n = 52     
       
3. 14 + x = 46; x = 32         YES  (14 + 32 = 46.  Correct !)
    
4. 17y = 85; y = 5             YES  (17*5 = 85.     Correct !)
        
5. 25 =   k _ 5   ; k = 5   
          
6. 2.5n = 45; n = 18           YES  (2.5*18 = 45.   Correct !)
   
7. 21 = m + 9; m = 11          NO   (11 + 9 = 20,   not 21)
     
8. 21 - h = 15; h = 6          YES  (21 - 6 = 15.   Correct !)
   
9. d - 4 = 19; d = 15          NO   (15 - 4 = 11,   not 19)   

10. 5 + x = 47; x = 52         NO   (5 + 52 = 57,   not 47)
   
11. w - 9 = 0; w = 9           YES  (9 - 9 = 0.     Correct !)

12. 5q = 31; q = 13            NO   (5*13 = 65,     not 31)
        
13. Each floor of a hotel has r rooms. On 8 floors, there are a total of 256 rooms. Write an equation to represent this situation,

                               8r = 256.

        H A P P Y     L E A R N I N G     to both of you !